OpenStudy (anonymous):

Miranda hiked to the top of a waterfall. She wanted to know the waterfall's height, but she was unable to measure it. Miranda created a proportion using the information from triangles PQS and TRS to calculate the height of the waterfall.

OpenStudy (anonymous):

How tall is the waterfall? (Round your answer to the nearest foot.)

OpenStudy (johnweldon1993):

As soon as we get the triangles information...we can help :P

OpenStudy (johnweldon1993):

Okay so we have the similar triangles...let the height of the waterfall = x \[\frac{ 730 }{ 480 } = \frac{ 730 + 256 }{ x }\] Also written as \[\frac{ 730 }{ 480 } = \frac{ 986 }{ x }\] Cross multiply and solve for 'x'

OpenStudy (anonymous):

730 ------ 473280

OpenStudy (johnweldon1993):

Other way around... 730 times x = 730x 986 times 480 = 473280 so we have 730x = 473280 now divide both sides by 730 730x = 473280 --------------- 730 730 x = ....?

OpenStudy (anonymous):

648.32

OpenStudy (johnweldon1993):

Right...and rounded to the nearest foot = ...?

OpenStudy (anonymous):

648

OpenStudy (johnweldon1993):

There you go! :)

OpenStudy (anonymous):

thanks :P

OpenStudy (johnweldon1993):

No problem *I'll gladly take your medal too >:)

OpenStudy (anonymous):

pfft please i owed that crap

OpenStudy (johnweldon1993):

haha I see how it is! GREEDY!!! :P

OpenStudy (anonymous):

hahahah no imma angel

OpenStudy (johnweldon1993):

Haha yeah I can see your halo XD NOT!

OpenStudy (anonymous):

no srsly ready O:)

OpenStudy (johnweldon1993):

Boo crappy Halo :P it's fake!!! haha ;P

OpenStudy (anonymous):

nah it aint its only fake with im pissed

OpenStudy (johnweldon1993):

hahaha then its >:) lol

OpenStudy (anonymous):

i aint mad yet

OpenStudy (johnweldon1993):

I can change that!!! :P

OpenStudy (anonymous):

nope u cant

OpenStudy (johnweldon1993):

You're right...I'm just too nice of a person...sigh...what a curse XD

OpenStudy (anonymous):

hhahahha i like nice people i hate fakes

OpenStudy (johnweldon1993):

Well then you should like me :)

OpenStudy (anonymous):

no i hate u

OpenStudy (johnweldon1993):

:O Halo = broken! xD